# OPRA-Vis: Visual Analytics System to Assist Organization-Public Relationship Assessment with Large Language Models

### v1

@@ III-A Organization-Public Relationship Assessment (OPRA)
The OPRA [7] is a framework devised by scholars in PR to evaluate relationships, featuring four fundamental concepts [13]: Trust, Satisfaction, Commitment, and Control Mutuality. Trust comprises competence, dependability, and integrity. Satisfaction relates to positive emotional responses. Commitment concerns long-term expectations, while Control Mutuality addresses perceptions of balanced control. OPRA provides metrics for gauging these relationships, offering valuable insight to maintain favorable public connections. These metrics are typically obtained via surveys, supporting relationship improvement strategies across sectors such as business and government, and helping organizations strengthen public engagement over both short- and long-term horizons. PR analysis involves two distinct classification tasks: the binary assessment of OPRA concepts (labeling opinions as True or False for each concept) and sentiment analysis, which classifies the emotional tone of public opinion as Positive or Negative.
@@ III-C Design Goals
Integrating PR theory with LLMs for labeling public opinion data faces three main challenges (C1–C3) as introduced in Section I. To address the challenge of insufficient training data for fine-tuning LLMs (G1), we implement the Gemma [17] model with CoT prompting, based on in-context learning. This method leverages textitFSEs provided by PR experts to label the data. Moreover, we have developed an LLM pipeline that represents OPRA-related clues in public opinion data as a continuous encoding (G2). Adopting a BERT [19] model reduces the dependence on human effort to classify sentiment by analyzing public opinion sentences. Sentiment classification enriches labeled data and broadens the analyses supported by our system (G2).
We combine visual analytics with CoT prompting that incorporates clues and reasons to audit and revise the model’s reasoning for labeling, supported by coordinated views and prompt editing (G3, G4, G5). CoT prompting elicits “think-aloud” rationales, making the LLM’s reasoning for labeling explicit and auditable (G3). Additionally, understanding the transformer’s attention mechanism is crucial for comprehending the LLM’s reasoning for labeling and for guiding prompt editing in CoT-based revisions. Therefore, attention visualization is devised to provide an attention mechanism for the transformer that influences LLM’s labeling (G3, G4). Moreover, public opinion data, i.e., freeform text posted on digital media, is diverse and noisy. Therefore, techniques that support usability are required for labeling, reasoning, and editing. We adopt sentence embeddings and dimensionality reduction to encode the certainty relevant to the OPRA concept (G2). This encoded certainty is then visualized in a scatter plot, with filtering applied based on certainty levels. The continuous encoding with certainty-based filtering for OPRA labeling improves usability and reduces human workload in visual analysis (G5).
@@ IV-A Concept Labeling with Clues and Reasons
Figure 2: The system overview of OPRA-Vis. OPRA-Vis consists of three tightly coordinated views. (a) is an OPRA Concept Space view to explore the certainty of concepts ($CoC$) relying on OPRA concepts using a scatter plot and provide the sentiment of words in sentences using tag clouds. (a-1) shows a distribution chart displaying data point density along the selected concept axis. (b) is a Data Management view. The Data Management view features a slider for data filtering, as presented in (b-1). The filter uses the CoCCoC range to update point visibility in the scatter plot in (a). The data table includes certainty, concept labels from LLM and experts, and data exclusion, as shown in (b-2). (c) is an LLM Reasoning & Editing view for analyzing LLM’s reasoning for labeling and editing prompts to adapt the user’s knowledge. This view provides an attention visualization for analyzing the effectiveness of sentences in prompting. (c-1) is a pop-up window that organizes the generated sentences, focus sentences, and attention scores. The pop-up window is activated by hovering over the dot representing the sentence of attention visualization. (c-2) shows buttons to open the prompt editor and run the assessment using the edited prompt.
@@ IV-C Certainty of Concepts (CoC)
Here, the max operation selects the highest attention score within each expert instruction, while the mean operation averages these maximum scores across multiple expert instructions for the same concept-label pair. The computed $CoC$ values are standardized and rescaled to $[0,1]$ for consistent comparison and intuitive interpretation: values near $1$ indicate high certainty for True OPRA concepts, values near 0 indicate high certainty for False OPRA concepts, and $0.5$ denotes neutral certainty. These $CoC$ values serve as quantitative indicators of concept certainty, supporting analytical evaluation.
@@ V-A OPRA Concept Space View
Figure 3: Position adjustment using the gravity model weighted by CoCCoC. (a) Before: Initial t-SNE layout of sentence embeddings forms dense clusters without clear concept alignment. (b) After: A gravity model uses True and False concept vertices ($\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ in Equation 2) as gravitational attractors, where each sentence is pulled toward its corresponding concept edge with force proportional to its certainty (CoCCoC). High-certainty sentences experience a stronger gravitational pull and move closer to concept edges. In contrast, low-certainty sentences experience weaker forces and remain near the center, resulting in eight radial structures separating concepts and certainty levels.
The OPRA Concept Space view, shown in Figure 2 (a), visualizes public opinion texts according to the four OPRA concepts—Trust, Satisfaction, Commitment, and Control Mutuality—using a scatter plot and tag clouds. We adopt an octagonal layout that uses the angle to indicate the direction of the concept-specific True or False opposing label and the radius to encode $CoC$ in Equation 1. Sentence embeddings are represented as points inside the octagon, while tag clouds are positioned around its perimeter. The embeddings are generated with $\text{BERT}_{\text{BASE}}$ [19], producing 768-dimensional vectors that are projected into 2D using t-SNE [48]. To better reflect OPRA-specific relationships, we further adjust the positions using Equation 2, which incorporates the $CoC$. The gravitational adjustment ensures that as $CoC$ increases, points move toward the corresponding concept edge. To achieve this adjustment, we employ a physics-inspired gravity model [49], where the vertices representing OPRA concepts act as gravitational attractors. Sentence positions are iteratively updated based on their $CoC$ values through a constrained simulation.
@@
Here, $\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ represent True and False concept vertices for concept $j$, $\mathbf{1}_{\{\cdot\}}$ is the indicator function, $\alpha_{\text{base}}$ is the base gravitational strength, $G$ is the gravitational constant, $r_{ij}=\|\mathbf{v}_{\text{target}}^{(i,j)}-\mathbf{p}_{i}^{(t)}\|_{2}$ is the Euclidean distance, $\mathbf{u}_{i}^{(t)}$ is the velocity vector, $\gamma$ is the damping factor, $\delta$ is the step size, $\epsilon_{1}$ prevents force divergence when $r_{ij}=0$, and $\epsilon_{2}$ prevents division by zero when computing directions. In our implementation, we use $\alpha_{\text{base}}=2.0$, $G=1.0$, $\gamma=0.8$, $\delta=0.1$, $\epsilon_{1}=0.01$, and $\epsilon_{2}=10^{-10}$. These parameters were empirically selected and can be adjusted. We will explore systematic optimization methods in future work. The gravitational force follows an inverse square law, where $\alpha_{ij}$ acts as the gravitational mass proportional to concept certainty. When $CoC_{ij}=0.5$, no gravitational force is applied as the sentence shows neutral certainty. Positions are constrained within a unit circle through normalization to maintain visual coherence. This iterative gravitational simulation continues for up to 200 iterations or until position changes fall below $10^{-4}$, creating spatial clustering where sentences are gravitationally attracted toward vertices corresponding to their respective concepts. Figure 3 illustrates this transformation: (a) shows the initial t-SNE layout, and (b) shows the adjusted plot reflecting $CoC$ values with gravitational attractors.
@@
The scatter plot enables interactive concept selection: clicking on an OPRA concept label on the concept edge updates the color encoding based on the $CoC$ values for that concept. We use a blue-yellow-red color scheme, where blue indicates low certainty (near 0), yellow represents neutral certainty (around 0.5), and red corresponds to high certainty (near 1). This three-color gradient aligns naturally with the $CoC$ scale, which is centered around the neutral value of 0.5. Figures 4 (a)–(d) illustrate scatter plots for the cases when the user selects Trust, Satisfaction, Commitment, and Control Mutuality, respectively.
To highlight regions of high and low certainty, we provide a distribution chart for the user-selected OPRA concept label from the scatter plot, as illustrated in Figure 2 (a-1). The data are aligned along an axis whose endpoints correspond to the concept’s True and False values. Each point is orthogonally projected onto this axis, producing a position in $[0,1]$ based on its $CoC$. The distribution is shown as a histogram with a $\log_{10}$-scaled y-axis to emphasize patterns across frequency ranges. Users can switch between linear, $\ln$, $\log_{2}$, and $\log_{10}$ scales by clicking the chart title, allowing flexible analysis of both sparse and dense regions.
@@ V-C LLM Reasoning & Editing View
Figure 6: LLM’s reasoning for labeling on Amazon product review data. (a) shows the low $CoC$s for Satisfaction in the scatter plot. (b) presents public opinion data where LLM and experts label differently. (c) reveals that LLM has a flawed reasoning for labeling. (d) is a prompt where the user revises the clues and displays the assessment results, which show that LLM’s reasoning for labeling has been revised.
@@ VI-A Linguistic Negation Correction
This use scenario involves 198 Amazon electronics reviews [51]. The OPRA concept labels as the ground-truth of these reviews are labeled by two PR researchers. To address LLM misclassifications, our system enables analyst-driven decision-making through prompt editing. This includes certainty-based filtering, sentiment-aware tag clouds, and attention visualization. Figure 6 provides an example of the relabeling process.
Low $CoC$ often arises from linguistic phenomena such as negation, introducing ambiguity in LLM’s assessments [53]. Figure 6 (a) illustrates this issue, where OPRA-Vis indicates low $CoC$s for Satisfaction. Using the lasso selection tool, we cluster the dense blue points. The Data Management view in (b) reveals that 41 sentences fall within the $\sim0$.3 CoCCoC range, and some points are mislabeled. (b) shows a mislabeled example, “The keyboard is easy to clean, […] the clear underneath”. While the expert labels this data point as True for Satisfaction, as indicated by “Label (Expert) = True”, the LLM labels it as False. This mismatch likely happens due to LLM’s limited interpretability of natural language.
@@
In Figure 6 (c), LLM misclassified the sentence “water doesn’t bother it too much”, which conveys a True meaning, as False. This is shown in CLUES (Sentence ID==9), where LLM focuses on the phrase “doesn’t bother it too much”, highlighting its difficulty in handling negation. Consequently, LLM inferred the SATISFACTION label (Sentence ID==11) as False, based on REASONING (Sentence ID==10). (c-1) visualizes that REASONING (Sentence ID==10) was most influenced by CLUES (Sentence ID==9), shown as an orange dot. For clarity, we excluded Sentence ID==0 and Sentence ID==10, both represented as red dots with the highest attention scores, from this inter-sentence attention analysis. Sentence ID==0 corresponds to Instructions and appears in nearly all generations, making it unsuitable for analyzing sentence-specific influence (see Sections V-C). Sentence ID==10 was also excluded as it is self-referential in the attention map. These findings guided prompt editing, as shown in Figure 7.
@@
Figure 6 (d) shows the updated model assessment following the edited prompt. The CLUES at Sentence ID=9 reflects the revised input, while the REASONING at Sentence ID=10 is newly generated based on these modified clues. Influenced by this reasoning, LLM corrects its assessment of Satisfaction to True. OPRA-Vis propagates the edit by updating the shared prompt template and the exemplar pool. Then it reassesses the entire dataset, updating labels, rationales, and $CoC$ values. This process clarifies how the prompt manages phenomena like negation, updates similar cases, increases $CoC$ near relevant concept edges, and reduces inconsistencies between the LLM and expert labels.
@@ VI-B Correcting Mislabeling Caused by Sentiment Bias
Figure 8 (a) displays the initial mislabeling. In a sentence from public opinion data indicating that the issue was resolved, the LLM focuses on positive-sentiment phrases, such as “fixed the problem”. It interprets these as evidence for the OPRA concept Trust and outputs True, while the expert label is False. This error reveals a reasoning shortcut: the LLM equates problem resolution or satisfaction with trustworthiness, rather than evaluating OPRA-informed clues such as transparency, fairness, and ethical conduct. The table emphasizes the discrepancy between “Label (LLM) = True” and “Label (Expert) = False”, along with low $CoC$, which aligns with sentiment-driven reasoning, as indicated by the orange-highlighted positive sentiment in the initial LLM reasoning. Figure 8 (b) shows the correction made after PR experts rewrite and expand the FSEs in the Prompt Editor to anchor better Trust to the OPRA-informed clues, as described in Section III-A. In the corrected reasoning, the LLM explicitly dismisses positive tone as sufficient evidence and instead grounds its assessment on trustworthiness in the REASONING text. This assessment aligns with the expert label of False, demonstrating that editing FSEs to align with concepts effectively mitigates sentiment bias and restores construct-centered labeling.

### v2

@@ III-A Organization-Public Relationship Assessment (OPRA)
The OPRA [7] is a framework devised by scholars in PR to evaluate relationships, featuring four fundamental concepts [13]: Trust, Satisfaction, Commitment, and Control Mutuality. Trust comprises competence, dependability, and integrity. Satisfaction relates to positive emotional responses. Commitment concerns long-term expectations, while Control Mutuality addresses perceptions of balanced control. OPRA provides metrics for gauging these relationships, offering valuable insight to maintain favorable public connections. These metrics are typically obtained via surveys, supporting relationship improvement strategies across sectors such as business and government, and helping organizations strengthen public engagement over both short- and long-term horizons. In PR analysis, OPRA concepts are evaluated at the sentence level. A sentence is labeled True for a concept if it contains identifiable clues supporting that concept; otherwise, it is labeled False. These binary labels denote the presence or absence of a concept, rather than the overall quality of the relationship. PR experts rely on cited clues and the distribution of True/False sentences to inform policy recommendations for strengthening organization–public relationships. Sentiment analysis complements OPRA labeling by capturing the emotional tone of a sentence (e.g., Positive or Negative), which is not conveyed by concept labels.
@@ III-C Design Goals
Integrating PR theory with LLMs for labeling public opinion data faces main challenges (C1–C3) as introduced in Section I. To address the challenge of insufficient training data for fine-tuning LLMs (G1), we implement the Gemma [17] with CoT prompting, based on in-context learning. This method leverages FSEs provided by PR experts to label the data. Moreover, we have developed an LLM pipeline that represents OPRA-related clues in public opinion data as a continuous encoding (G2). Adopting a BERT [19] reduces the dependence on human effort to classify sentiment by analyzing public opinion sentences. Sentiment classification enriches labeled data and broadens the analyses supported by our system (G2).
We combine VA with CoT prompting that incorporates clues and reasons to audit and revise the model’s reasoning for labeling, supported by coordinated views and prompt editing (G3, G4, G5). CoT prompting elicits “think-aloud” rationales, making the LLM’s reasoning for labeling explicit and auditable (G3). Additionally, understanding the transformer’s attention mechanism is crucial for comprehending the LLM’s reasoning for labeling and for guiding prompt editing in CoT-based revisions. Therefore, attention visualization is devised to provide an attention mechanism for the transformer that influences LLM’s labeling (G3, G4). Moreover, public opinion data, i.e., freeform text posted on digital media, is diverse and noisy. Therefore, techniques that support usability are required for labeling, reasoning, and editing. We adopt sentence embeddings and dimensionality reduction to encode the certainty relevant to the OPRA concept (G2). This encoded certainty is then visualized in a scatter plot, with filtering applied based on certainty levels. The continuous encoding with certainty-based filtering for OPRA labeling improves usability and reduces human workload in visual analysis (G5).
@@ IV-A Concept Labeling with Clues and Reasons
Figure 2: The system overview of OPRA-Vis. OPRA-Vis consists of three tightly coordinated views. (a) is an OPRA Concept Space view to explore the certainty of concepts ($CoC$) relying on OPRA concepts using a scatter plot and provide the sentiment of words in sentences using tag clouds. (a-1) shows a distribution chart displaying data point density along the selected concept axis. (b) is a Data Management view. The Data Management view features a slider for data filtering, as presented in (b-1). The filter uses the $CoC$ range to update point visibility in the scatter plot in (a). The data table includes certainty, concept labels from LLM and experts, and data exclusion, as shown in (b-2). (c) is an LLM Reasoning & Editing view for analyzing LLM’s reasoning for labeling and editing prompts to adapt the user’s knowledge. This view provides an attention visualization for analyzing the effectiveness of sentences in prompting. (c-1) is a pop-up window that organizes the generated sentences, focus sentences, and attention scores. The pop-up window is activated by hovering over the dot representing the sentence of attention visualization. (c-2) shows buttons to open the prompt editor and run the assessment using the edited prompt.
@@ IV-C Certainty of Concepts (CoC)
Here, the max operation selects the highest attention score within each expert instruction, while the mean operation averages these maximum scores across multiple expert instructions for the same concept-label pair. The computed $CoC$ values are standardized and rescaled to $[0,1]$ for consistent comparison and intuitive interpretation. These $CoC$ values serve as quantitative indicators of concept certainty, supporting analytical evaluation.
@@ V-A OPRA Concept Space View
Figure 3: Position adjustment using the gravity model weighted by CoCCoC. (a) Before: Initial t-SNE layout of sentence embeddings forms dense clusters without clear concept alignment. (b) After: A gravity model uses True and False label vertices ($\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ in Equation 2) as gravitational attractors, where each sentence is pulled toward its corresponding concept edge with force proportional to its CoCCoC deviation from 0.5. Sentences with CoCCoC values distant from 0.5 are pulled more strongly toward concept edges, while those near 0.5 experience weaker forces and remain closer to the center. This produces eight radial structures that separate concepts by certainty levels.
The OPRA Concept Space view, shown in Figure 2 (a), visualizes public opinion texts according to the four OPRA concepts—Trust, Satisfaction, Commitment, and Control Mutuality—using a scatter plot and tag clouds. We adopt an octagonal layout that uses the angle to indicate the direction of the concept-specific True or False opposing label and the radius to encode $CoC$ in Equation 1. Sentence embeddings are represented as points inside the octagon, while tag clouds are positioned around its perimeter. The embeddings are generated with $\text{BERT}_{\text{BASE}}$ [19], producing 768-dimensional vectors that are projected into 2D using t-SNE [48]. To better reflect OPRA-specific relationships, we further adjust the positions using Equation 2, which incorporates the $CoC$. We employ a physics-inspired gravity model [49], where the vertices representing OPRA concepts act as gravitational attractors. Sentence positions are iteratively updated based on their $CoC$ values through a constrained simulation.
@@
Here, $\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ represent True and False label vertices for concept $j$, $\mathbf{1}_{\{\cdot\}}$ is the indicator function, $\alpha_{\text{base}}$ is the base gravitational strength, $G$ is the gravitational constant, $r_{ij}=\|\mathbf{v}_{\text{target}}^{(i,j)}-\mathbf{p}_{i}^{(t)}\|_{2}$ is the Euclidean distance, $\mathbf{u}_{i}^{(t)}$ is the velocity vector, $\gamma$ is the damping factor, $\delta$ is the step size, $\epsilon_{1}$ prevents force divergence when $r_{ij}=0$, and $\epsilon_{2}$ prevents division by zero when computing directions. In our implementation, we use $\alpha_{\text{base}}=2.0$, $G=1.0$, $\gamma=0.8$, $\delta=0.1$, $\epsilon_{1}=0.01$, and $\epsilon_{2}=10^{-10}$. These parameters were empirically selected and can be adjusted. We will explore systematic optimization methods in future work. The gravitational force follows an inverse square law, where $\alpha_{ij}$ acts as the gravitational mass proportional to the deviation of $CoC_{ij}$ from 0.5. When $CoC_{ij}~=~0.5$, no gravitational force is applied as the sentence shows intermediate $CoC$ values. Positions are constrained within a unit circle through normalization to maintain visual coherence. This iterative gravitational simulation continues for up to 200 iterations or until position changes fall below $10^{-4}$, creating spatial clustering where sentences are gravitationally attracted toward vertices corresponding to their respective concept labels. Figure 3 illustrates this transformation: (a) shows the initial t-SNE layout, and (b) shows the adjusted plot reflecting $CoC$ values with gravitational attractors.
@@
The scatter plot enables interactive concept selection. Clicking on an OPRA concept label on the concept edge updates the color encoding based on the CoC values for that concept using a blue-yellow-red color scheme. Figures 4 (a)–(d) illustrate scatter plots for the cases when the user selects Trust, Satisfaction, Commitment, and Control Mutuality, respectively.
To illustrate how $CoC$ values translate into positions along the concept axis, we provide a distribution chart for the user-selected OPRA concept label from the scatter plot, as illustrated in Figure 2 (a-1). The axis spans from False to True, and each data point is orthogonally projected onto the axis, yielding a position within $[0,1]$. The distribution is shown as a histogram with bars stacked by $CoC$ values and a $\log_{10}$-scaled y-axis to emphasize patterns across frequency ranges. Users can toggle between linear, $\ln$, $\log_{2}$, and $\log_{10}$ scales by clicking on the chart title, enabling flexible analysis of both sparse and dense regions of the position distribution.
@@ V-C LLM Reasoning & Editing View
Figure 6: LLM’s reasoning for labeling on Amazon product review data. (a) shows the OPRA Scpace view with Satisfaction selected by the user. b) presents public opinion data where LLM and experts label differently. (c) reveals that LLM has a flawed reasoning for labeling. (d) is a prompt where the user revises the clues and displays the assessment results, which show that LLM’s reasoning for labeling has been revised.
@@ VI-A Linguistic Negation Correction
This use scenario involves 198 Amazon product reviews [51]. The OPRA concept labels as the ground-truth of these reviews are labeled by two PR researchers. To address LLM misclassifications, our system enables analyst-driven decision-making through prompt editing. This includes certainty-based filtering, sentiment-aware tag clouds, and attention visualization. Figure 6 provides an example of the relabeling process.
Guided by the distribution chart in Figure 6 (a), we partition the data into three $CoC$ groups, as shown within the three gray dotted frames. While exploring the group (a-2) with $CoC$ in the range of 0 to 0.3 using the filtering on the Data Management view in (b), we capture that 41 sentences surface several mismatches between expert and LLM labels. (b) shows a representative case, “The keyboard is easy to clean, […] the clear underneath”. The expert labels this sentence True for Satisfaction, as indicated by “Label (Expert) = True”, whereas the LLM labels it False. As we examine the LLM’s reasoning for labeling in (c), we find that linguistic negation underlies this error. Negation is a well-known source of ambiguity in LLM’s labeling tasks [53].
@@
In Figure 6 (c), LLM misclassified the sentence “water doesn’t bother it too much”, which conveys a True meaning, as False. This is shown in CLUES (Sentence ID==9), where LLM focuses on the phrase “doesn’t bother it too much”, highlighting its difficulty in handling negation. Consequently, LLM inferred the SATISFACTION label (Sentence ID==11) as False, based on REASONING (Sentence ID==10). (c-1) visualizes that REASONING (Sentence ID==10) was most influenced by CLUES (Sentence ID==9), shown as an orange dot. For clarity, we excluded Sentence ID==0 and Sentence ID==10, both represented as red dots with the highest attention scores, from this inter-sentence attention analysis. Sentence ID==0 corresponds to Instructions and appears in nearly all generations, making it unsuitable for analyzing sentence-specific influence (see Section V-C). Sentence ID==10 was also excluded as it is self-referential in the attention map. These findings guided prompt editing, as shown in Figure 7.
@@
Figure 6 (d) shows the updated model assessment following the edited prompt. The CLUES at Sentence ID=9 reflects the revised input, while the REASONING at Sentence ID=10 is newly generated based on these modified clues. Influenced by this reasoning, LLM corrects its assessment of Satisfaction to True. OPRA-Vis propagates the edit by updating the shared prompt template and the exemplar pool. Then it reassesses the entire dataset, updating labels and rationales. This process clarifies how the prompt manages phenomena like negation, updates similar cases, and reduces inconsistencies between the LLM and expert labels.
@@ VI-B Correcting Mislabeling Caused by Sentiment Bias
Figure 8 (a) displays the initial mislabeling. In a sentence from public opinion data indicating that the issue was resolved, the LLM focuses on positive-sentiment phrases, such as “fixed the problem”. It interprets these as evidence for the OPRA concept Trust and outputs True, while the expert label is False. This error reveals a reasoning shortcut: the LLM equates problem resolution or satisfaction with trustworthiness, rather than evaluating OPRA-informed clues such as transparency, fairness, and ethical conduct. The table emphasizes the discrepancy between “Label (LLM) = True” and “Label (Expert) = False”, which aligns with sentiment-driven reasoning, as indicated by the orange-highlighted positive sentiment in the initial LLM reasoning. Figure 8 (b) shows the correction made after PR experts rewrite and expand the FSEs in the Prompt Editor to anchor better Trust to the OPRA-informed clues, as described in Section III-A. In the corrected reasoning, the LLM explicitly dismisses positive tone as sufficient evidence and instead grounds its assessment on trustworthiness in the REASONING text. This assessment aligns with the expert label of False, demonstrating that editing FSEs to align with concepts effectively mitigates sentiment bias and restores construct-centered labeling.